AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 300     Entry time: Sun Feb 27 09:00:09 2022
Author: RSS TD 
Subject: Current status 

10:00: 95K-->120K

10:08: 120K --> 95K

Ring current ~7e7 particles/s

Merger, Tape server (WS3): 1414 kB/s writing ok

Leakage current (WS4): 3.43 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 10-xy-2022-02-27_10-02-27.png  273 kB  | Hide | Hide all
10-xy-2022-02-27_10-02-27.png
Attachment 2: 10-ey-2022-02-27_10-02-04.png  167 kB  | Hide | Hide all
10-ey-2022-02-27_10-02-04.png
Attachment 3: 10-ex-2022-02-27_10-01-46.png  166 kB  | Hide | Hide all
10-ex-2022-02-27_10-01-46.png
Attachment 4: 10-ey-rate-2022-02-27_10-09-54.png  156 kB  | Hide | Hide all
10-ey-rate-2022-02-27_10-09-54.png
Attachment 5: 10-ex-rate-2022-02-27_10-09-33.png  156 kB  | Hide | Hide all
10-ex-rate-2022-02-27_10-09-33.png
Attachment 6: 10-ex-rate-95k-2022-02-27_10-13-33.png  154 kB  | Hide | Hide all
10-ex-rate-95k-2022-02-27_10-13-33.png
Attachment 7: 10-ey-rate-95K-2022-02-27_10-17-09.png  156 kB  | Hide | Hide all
10-ey-rate-95K-2022-02-27_10-17-09.png
ELOG V3.1.4-unknown